OpenStudy (anonymous):

f(x)=3x^2-9x+2)(-x^2+3)

OpenStudy (anonymous):

@TheThunderChimp ??

OpenStudy (anonymous):

Do you have to evaluate that at a point or expand it out?

OpenStudy (anonymous):

i honeslty dont know my teacher never meantioned anything like that

OpenStudy (anonymous):

Well that looks more like an answer than a question! You can rewrite it (3x^2)*(-x^2)+(3x^2)*3-9x*(-2x)-9x*(3)+2*(-x^2)+2*3

OpenStudy (anonymous):

oooohhh i forgot you gotta find g(-6)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

You mean f(-6)? That would be (3*(-6)^2-9*(-6)+2)*((-6)^2+3)=164*39=6396, I think